# CLASSMix: Adaptive stain separation-based contrastive learning with pseudo labeling for histopathological image classification

### v1

@@ 3.2 Adaptive stain separation
Stain separation is introduced to separate different types of stains present in histological images [41, 49, 57], such as separating Hematoxylin and Eosin(H&E) images into Hematoxylin images and Eosin images. The process has many challenges due to the variation of stains caused by different manufacturers, storage conditions and staining procedures. In our model, a simple and unsupervised method for adaptive stain separation [37] is applied. The main idea is to first project all pixels from RGB space to Optical Density (OD) space, where stain components are formed by linear combinations. Inside the OD space, the appropriate Hematoxylin and Eosin vectors are calculated. Then all the RGB pixels can be projected to the Hematoxyling portion and Eosin portion to form Hematoxylin images and Eosin images. The images are further normalized to mitigate the effect of variations on the brightness and dosage of stains. More detailed algorithm can be found in Appendix A.
3.3 Augmentations on original RGB images, Hematoxylin images and Eosin images
We perform color-jittering augmentations on brightness, contrast and saturation of original images, followed by other image augmentations on Hematoxylin images and Eosin images, including random crops, random rotations and random flips. Independent jittering of brightness on Hematoxylin images and Eosin images is also applied. The extra augmentations on original images provide more diversities in the training samples and help generalize more robust features.
@@ 3.4 Contrastive learning
The inputs with two different views are usually adopted in co-training. The general presumptions of co-training are that each input view carries sufficient information for the task and the views should be more independent of each other to achieve good results. Our preliminary work [54] has demonstrated that Hematoxylin and Eosin channels fulfill those requirements perfectly in contrastive co-training, especially compared to using the Red, Green or Blue channels in co-training.
@@
where $x_{k}$ represents another random sample, $\parallel a\parallel_{2}$ denotes the L2 norm of vector $a$, and $m$ serves as a margin hyperparameter. This contrastive loss term forms a positive pair whose $f_{H}(x)$ and $f_{E}(x)$ originate from the same sample and a negative pair whose $f_{H}(x)$ and $f_{E}(x)$ come from different samples. After training, the two features from the same samples tend to get closer in the shared latent feature space.
@@ 3.5 MixUp
MixUp [56] is applied on both labeled and unlabeled samples. First, we assign pseudo labels to unlabeled samples. Assume $x_{i}$ is an original image sample from unlabeled training set $U$, we augment $x_{i}$ $K$ times and get the average of softmax predictions from the model. The result is sharpened later to lower the entropy of label distribution. The formulas are as follows:
@@
$x_{i,k}$ means the $k$-th augmentation of sample $x_{i}$, $T$ is a temperature hyperparameter to control the sharpness, $y_{(c)}$ means the $c$-th element in one-hot label encoding $y$, $C$ means the total number of classes, and $y_{i}$ is the virtual one-hot label encoding of sample $x_{i}$.
@@
After MixUp, we define $x^{\prime}$ as labeled sample if $x_{i}$ is from labeled set to form new labeled set $L^{\prime}$, the remaining samples form new unlabeled set $U^{\prime}$ with virtual labels. Since the $\lambda^{\prime}$ is a random number during training, there would be a lot more virtual samples compared to the original samples.
@@ 3.6 Loss function
The total loss is formed by the summation of losses from label prediction and the contrastive learning. Inspired by MixMatch [3], we use cross-entropy loss in labeled set $L^{\prime}$ and squared $L_{2}$ loss between predictions and virtual labels in unlabeled set $U^{\prime}$ considering squared $L_{2}$ loss is less sensitive to incorrect predictions. The total loss in a batch can be written as
@@
$y_{i}$ is one-hot encoding of labels (in $L^{\prime}$) or virtual labels (in $U^{\prime}$), $\hat{y}_{j}$ is the label prediction from the model, C is the number of classes, $|L^{\prime}|$ and $|U^{\prime}|$ means the size of labeled set and unlabeled set in a batch. $\lambda_{U^{\prime}}$ and $\lambda_{C}$ are hyperparameters that control the weights of squared $L_{2}$ loss and contrastive learning loss.
@@ 4.1 Datasets
To evaluate our method, we applied our CLASSMix model on Utah clear cell renal cell carcinoma (ccRCC) dataset and The Cancer Genome Atlas Program (TCGA) dataset seperately, and compared results with other semi-supervised and self-supervised classification methods.
@@
In the Utah ccRCC dataset, there are 49 Whole Slide Images (WSIs) from 49 patients. First, a pathologist drew polygons inside WSIs to mark areas with certain growth patterns, which were subsequently verified by another pathologist. We randomly split the WSIs into 32, 10, and 7 WSIs for training, validation and test set. The WSIs were then cropped into 400$\times$400 tiles in 10X resolution with sampling step set to 200 pixels inside each labeled polygon. The same process was applied to crop tiles outside polygons in 32 training WSIs to collect unlabeled tiles to form a semi-supervised learning task for a 4-class classification. In detail, there are 28497, 2044, 2522, 4115 tiles in the category of Normal tissue, Low risk cancer, High risk cancer, and Necrosis, respectively, in the labeled training set extracted from polygons in 32 WSIs. Additionally, there are 171,113 unlabeled training tiles, 5472, 416, 334, 2495 tiles for each category in validation set, and 7263, 598, 389, 924 tiles for each category in test set. Tiles with predominantly background areas were removed, along with those tainted by ink. Tile examples can be seen in Figure 2(a).  [^1] The Utah ccRCC dataset is available through a transfer agreement by contacting authors.
@@
For the TCGA ccRCC dataset, we have in total 420 WSIs from 420 patients. 150 of them were labeled by a pathologist through drawing polygons with annotations and verified by another pathologist, and the remaining 270 WSIs were unlabeled. If pathologists had disagreements or concerns about labels of polygons, then those annotations were abandoned. The resolution we trained on was 20X. To make it a more challenging task and show the effectiveness of semi-supervised learning, we split the 150 WSIs into 30, 60 and 60 WSIs for labeled training, validation and test set. The tile cropping process was the same as in the Utah ccRCC dataset except for different sampling steps. We chose 200 pixels as sampling step for labeled training set, but 400 pixels for validation and test set, considering they contain enough tiles. We cropped foreground tiles outside polygons in 30 labeled training WSIs, as well as foreground tiles across the 270 unlabeled WSIs to form unlabeled samples by setting sampling step to 400 pixels. The tiles maintained 400$\times$400 size at all times. We split the labeled tiles into 3 categories: Normal tissue, Cancer and Necrosis to perform 3 class semi-supervised learning classification tasks. Tile examples can be seen in Figure 2(b). In summary, we have 84578, 180471, 7932 labeled training tiles, respectively, in the category of Normal tissue, Cancer and Necrosis, as well as 19638, 79382, 1301 validation tiles, and 15323, 62565, 6168 test tiles for each category. The number of unlabeled training tiles is 1,373,684. The TCGA ccRCC dataset is a publicly available dataset, and our annotations for each WSI are available by contacting authors.
4.2 Adaptive stain separation
In experiments, we maintained slide-by-slide level adaptive stain separation based on the following considerations: The stain styles are various across whole slide images due to different conditions. But within one whole slide image, most conditions, such as stain manufacturer, storage condition, and staining procedure, remain consistent. On the other hand, if adaptive stain separation is applied patch-by-patch, there are not enough pixel samples inside each patch to obtain robust stain separation results. For example, in Figure 2(a), the top image in Necrosis category has very limited Hematoxylin portion, which makes stain separation less trustful at patch-by-patch level.
The RGB to H/E conversion matrices for each WSI were already calculated before the CLASSMix model training. We provide our original code for adaptive stain separation based on algorithms in [37]. Please contact authors to get access to the code.
@@ 4.2 Adaptive stain separation
Table 1: Performance of different classification models on Utah ccRCC dataset and TCGA ccRCC dataset. Mean accuracy and standard deviation on test set are calculated. The encoders of self-supervised learning foundational models including Barlow Twins, SwAV, MoCov2, and ViT-DINO are pretrained on large amount of histopathological images by [29] and then frozen. The results with the best mean accuracy are shown in bold.
@@
Fully-supervised ResNet $88.85\pm 2.66\%$ $72.11\pm 0.41\%$
@@
Transfer learning ResNet (Freeze) $85.68\pm 0.18\%$ $73.17\pm 0.42\%$
ViT (Freeze) $82.51\pm 0.15\%$ $73.17\pm 0.46\%$
Self-supervised BT $90.64\pm 1.21\%$ $80.81\pm 1.79\%$
SwAV $91.22\pm 0.67\%$ $77.23\pm 3.61\%$
MoCoV2 $87.35\pm 0.21\%$ $80.54\pm 1.75\%$
ViT-DINO $92.11\pm 0.86\%$ $77.61\pm 3.64\%$
Semi-supervised FixMatch $91.58\pm 0.65\%$ $83.34\pm 2.53\%$
@@
CLASSMix $\mathbf{95.35\pm 0.46\%}$ $\mathbf{92.13\pm 0.89\%}$
@@ 4.3 Experiment settings
For a fair comparison, we used ImageNet pretrained ResNet18 model [23] as backbone for all Convolutional Neural Network (CNN) models in semi-supervised learning. We compared our CLASSMix model with other state-of-the-art semi-supervised learning models, including FixMatch [47] and MixMatch [3]. In CLASSMix model training, we performed jittering of brightness, contrast and saturation on original RGB images, as well as image augmentations on H images and E images after stain separation, such as random rotation, random crop to $256\times 256$, random flip and jittering of brightness. The image augmentations on H images and E images are independent, for example, they could have different rotation angles in random rotation. Since H images and E images are one-channel images, we summed the pretrained weights on the first layer of ResNet18 as initialization weights.
@@
We used Root Mean Squared Propagation optimization method with a decaying learning rate. A batch size of 64 was chosen for experiments. Balanced sampler was used to address the imbalance in the number of labels in the training set. In experiments with the Utah ccRCC dataset, we chose to have 32 labeled samples in each batch, with 8 samples for each category, while the remaining 32 samples were unlabeled. In experiments with the TCGA ccRCC dataset, we chose to have 33 labeled samples in each batch, with 11 samples for each category, and the remaining 31 samples would be unlabeled. In validation and test, we first calculated classification accuracies for each category, and then averaged them to get balanced validation accuracy and test accuracy. This approach ensures that each category holds equal importance, irrespective of the actual number of tiles for each category. The hyperparameters were fine-tuned for all models to get the best validation accuracy, then the test accuracy on that hyperparameter setting was the final performance. We found that most hyperparameters didn’t need to be changed between Utah ccRCC dataset and TCGA ccRCC dataset. We chose decaying learning rate with initial value set to $10^{-4}$. We set number of augmentation times K to 2, temperature $T$ to 0.5, $\alpha$ to 2, margin $m$ in contrastive loss to 37 and unlabeled $L_{2}$ loss weight to 7.5 for both datasets. We set contrastive loss weight to 0.1 for experiments on the Utah ccRCC dataset, and contrastive loss weight to 0.001 for the TCGA ccRCC dataset. For each epoch, we ran 1000 iterations and checked validation accuracy once. The total epochs were large enough to ensure convergence. The training would stop if the best validation accuracy was no longer updated for more than 100 epochs. Each experiment was repeated three times to obtain the average test accuracy and standard deviation.
@@
The experiment platform is Python 3.7.11, Pytorch 1.9.0, torchvision 0.10.0, and CUDA 10.2. The GPUs we used are NVIDIA TITAN RTX. During CLASSMix model training, it usually took 13 GB GPU memory and 12 minutes for each training epoch. The code of our CLASSMix model is publicly available, please contact authors for the access to the code.
4.4 Results
We tested our CLASSMix model along with other state-of-the-art semi-supervised models, including FixMatch and MixMatch by conducting semi-supervised ccRCC histopathological image classification tasks on Utah ccRCC dataset and TCGA ccRCC dataset. The CLASSMix without pseudo labeling using MixUp was tested as a reference, and we call it CLASS model. The baseline ResNet18 [23] and Vision Transformer (ViT) [14] were also tested as references, where only labeled samples were used in training. The ResNet18 and ViT were pre-trained on ImageNet and then trained with two options, train all the parameters (fully-supervised learning), or freeze the encoder part and only train the final layer (transfer learning). All experiments were run three times to obtain mean test accuracy and standard deviation. Each test accuracy was calculated based on the mean of each category’s test accuracy to address the unbalanced nature of the numbers of each category in the test set.
As shown in Table 1, our CLASSMix model outperforms other state-of-the-art models by a large margin on both datasets. We achieved a test accuracy of $95.35\pm 0.46\%$ on Utah ccRCC dataset, and $92.13\pm 0.89\%$ on TCGA ccRCC dataset. The ResNet18 results from Table 1 indicates that it is a more challenging task on TCGA ccRCC dataset, whose test accuracy dropped by more than $15\%$. One possible reason is that the samples from Utah ccRCC dataset have less variation. For instance, the slides follow the same staining procedure, and the storage conditions are uniform. However, on TCGA ccRCC dataset, slides come from different institutions, leading to more variation in the quality of slides. Additionally, the Necrosis tiles on TCGA training set only come from 3 WSIs, which causes more difficulties in learning robust Necrosis features. As shown in the recall section of Table 4 in Appenix B, the pseudo labeling with MixUp raised the Necrosis’s test accuracy from around 60% to around 87% on TCGA ccRCC dataset, demonstrating that MixUp is especially effective on classes with very limited number of samples. In conclusion, our CLASSMix model achieves top results and proves to be the best on a variety of tasks. We also provide recall, precision and F-score for each run, with details available in Appendix B.
@@ 4.4 Results
Table 2: Ablation studies of CLASSMix models on Utah ccRCC dataset and TCGA ccRCC dataset. Mean accuracy and standard deviation on test set are calculated. The results with best mean accuracy are shown in bold.
@@
CLASSMix $\mathbf{95.35\pm 0.46\%}$ $\mathbf{92.13\pm 0.89\%}$
@@
CLASS without adaptive stain separation (use fixed stain seperation) $92.43\pm 0.34\%$ $79.89\pm 0.69\%$
@@
4.5 Quantitative comparison to Self-supervised learning foundation models
To demonstrate our superior performance in histopathology image classification by adopting adaptive stain separation-based contrastive learning and pseudo labeling with MixUp, we did extra self-supervised learning experiments for comparison. We employed a fully-connected classification network built upon pre-trained models from [29], including Barlow Twins [53], SwAV [6], MoCov2 [12], and ViT-DINO [7]. The encoder part was frozen after pre-training on a large amount of histopathological images by [29], and the final fully-connected classification layer was trained on our labeled training samples. As shown in Table 1, our proposed models, CLASS and CLASSMix outperformed those self-supervised learning models on ccRCC datasets even though they were pre-trained on very large histological images. The recall, precision and F-score of those self-supervised learning models can also be found in Appendix B.

### v4

@@ 3.2 Adaptive stain separation
Stain separation is introduced to separate different types of stains present in histological images  [44, 52, 61], such as separating Hematoxylin and Eosin (H&E) images into Hematoxylin images and Eosin images. The process has many challenges due to the variation of stains caused by different manufacturers, storage conditions and staining procedures. In our model, a simple and unsupervised method for adaptive stain separation  [40] is applied. The main idea is to first project all pixels from RGB space to Optical Density (OD) space, where stain components are formed by linear combinations. Inside the OD space, the appropriate Hematoxylin and Eosin vectors are calculated. Then all the RGB pixels can be projected to the Hematoxyling portion and Eosin portion to form Hematoxylin images and Eosin images. The images are further normalized to mitigate the effect of variations on the brightness and dosage of stains. A more detailed description of the algorithm can be found in Appendix A.
Figure 2: Contrastive learning on Hematoxyling images and Eosin images
3.3 Augmentations on original RGB, Hematoxylin and Eosin images
We perform color-jittering augmentations on brightness, contrast and saturation of original images, followed by image augmentations on Hematoxylin images and Eosin images, including random crops, random rotations and random flips. Independent jittering of brightness on Hematoxylin images and Eosin images is also applied. The extra augmentations on the original RGB images provide increased diversity in the training samples and help learn more general features.
@@ 3.4 Contrastive learning
Inputs from two different views are usually adopted in co-training. The general presumptions of co-training are that each input view carries sufficient information for the task and the views should be highly independent of each other to achieve good results. Our preliminary work  [58] has demonstrated that Hematoxylin and Eosin channels fulfill these requirements in contrastive co-training.
@@
where $x_{k}$ represents another random sample, $\parallel a\parallel_{2}$ denotes the L2 norm of vector $a$, and $m$ serves as a margin hyperparameter. This contrastive loss term forms a positive pair whose $f_{H}(x)$ and $f_{E}(x)$ originate from the same sample and a negative pair whose $f_{H}(x)$ and $f_{E}(x)$ come from different samples. Figure 2 shows an example of how the positive pairs and the negative pairs are formed.
During training, the H and E features from the same sample are pulled closer in the shared latent feature space, while the features from different samples are pushed further away.
@@ 3.5 MixUp
MixUp  [60] is applied on both labeled and unlabeled samples. First, we assign pseudo-labels to unlabeled samples. Assuming $x_{i}$ is an original unlabeled sample from unlabeled training set $U$, we augment it $K$ times and get the average of softmax predictions from the model. The result is sharpened to lower the entropy of label distribution using
@@
where $x_{i,k}$ iz the $k$-th augmentation of sample $x_{i}$, $T$ is a temperature hyperparameter to control the sharpness, $y_{(c)}$ is the $c$-th element in one-hot label encoding $y$, $C$ is the number of classes, and $y_{i}$ is the virtual one-hot label encoding of sample $x_{i}$.
@@
After MixUp, we define $x^{\prime}$ as labeled sample if $x_{i}$ is from labeled set to form new labeled set $L^{\prime}$, the remaining samples form new unlabeled set $U^{\prime}$ with virtual labels. Since $\lambda^{\prime}$ is random, sample diversity is greatly increased.
@@ 3.6 Loss function
The total loss is formed by the summation of losses from label prediction and contrastive learning. Inspired by MixMatch  [3], we use cross-entropy loss in labeled set $L^{\prime}$ and squared $L_{2}$ loss between predictions and virtual labels in unlabeled set $U^{\prime}$ considering squared $L_{2}$ loss is less sensitive to incorrect predictions. The total batch loss can be written as
@@
$y_{i}$ is one-hot encoding of labels (in $L^{\prime}$) or virtual labels (in $U^{\prime}$), $\hat{y}_{j}$ is the label prediction from the model, C is the number of classes, $|L^{\prime}|$ and $|U^{\prime}|$ are the size of labeled set and unlabeled set in a batch. $\lambda_{U^{\prime}}$ and $\lambda_{C}$ are hyperparameters that control the weights of squared $L_{2}$ loss and contrastive learning loss.
@@ 4.1 Datasets
To evaluate our method, we applied CLASS-M on the Utah ccRCC dataset and TCGA ccRCC dataset separately, and compared results with other semi-supervised and self-supervised classification methods.
@@
In the Utah ccRCC dataset, there are 49 WSIs from 49 patients. First, a pathologist drew polygons inside WSIs to mark areas with certain growth pattern labels, which were subsequently verified by another pathologist. We randomly split the WSIs into 32, 10, and 7 WSIs for training, validation and test set. The WSIs were then cropped into 400$\times$400 tiles at 10X resolution with stride set to 200 pixels inside each labeled polygon. The same process was applied to crop tiles outside polygons in 32 training WSIs to collect unlabeled tiles to form a semi-supervised learning task for a 4-class classification. In detail, there are 28497, 2044, 2522, 4115 tiles in the category of Normal/Benign, Low risk cancer, High risk cancer, and Necrosis, respectively, in the labeled training set extracted from polygons in 32 WSIs. Additionally, there are 171,113 unlabeled training tiles, 5472, 416, 334, 2495 tiles respectively for each category in the validation set, and 7263, 598, 389, 924 tiles respectively for each category in the test set. Tiles with predominantly background areas were removed, along with those tainted by ink. Tile examples can be seen in Figure 3(a).  [^1] The Utah ccRCC dataset is available through a transfer agreement by contacting the authors.
@@
For the TCGA ccRCC dataset, we have in total 420 WSIs from 420 patients. 150 of them were labeled by a pathologist through drawing polygons with annotations and verified by a second pathologist. The remaining 270 WSIs were treated unlabeled. If pathologists had disagreements or concerns about labels of polygons, then those annotations were abandoned. The resolution we trained on is 20X. To make it a more challenging task and show the effectiveness of semi-supervised learning, we split the 150 labeled WSIs into 30, 60 and 60 WSIs for training, validation and test set. The tile cropping process was the same as in the Utah ccRCC dataset except for different strides. We chose 200 pixels as stride for labeled training set, but 400 pixels for validation and test set, considering they contain a large number of tiles. We cropped foreground tiles outside polygons in the 30 labeled training WSIs, as well as foreground tiles across the 270 unlabeled WSIs to form unlabeled training samples by setting cropping stride to 400 pixels. The cropped tiles maintained 400$\times$400 size at all times. We split the labeled tiles into 3 categories: Normal/Benign, Cancer and Necrosis to perform 3-class semi-supervised classification tasks. Tile examples can be seen in Figure 3(b). In summary, there are 84578, 180471, 7932 labeled training tiles, respectively, in the category of Normal/Benign, Cancer and Necrosis, as well as 19638, 79382, 1301 validation tiles, and 15323, 62565, 6168 test tiles respectively for each category. The number of unlabeled training tiles is 1,373,684. The TCGA ccRCC dataset is a publicly available dataset, and our annotations for the 150 WSIs are accessible at github.com/BzhangURU/Paper_CLASS-M/tree/main/Section1_get_tiles_from_WSIs, where we also provide our code to perform tile cropping. The tiles described above can be easily generated through our code by reading WSIs and annotations as inputs.
@@
Table 1: Performance of different classification models on Utah ccRCC dataset and TCGA ccRCC dataset. Mean accuracy and standard deviation on test set are calculated. The supervised setting was only trained on labeled samples. The encoders of self-supervised learning models including Barlow Twins, SwAV, MoCo v3, and ViT-DINO are pre-trained on both labeled and unlabeled samples. The results with the best mean accuracy are shown in bold.
@@
Supervised (labeled images only) ResNet $88.85\pm 2.66\%$ $72.11\pm 0.41\%$
@@
Self-supervised (pre-trained on labeled and unlabeled images) Barlow Twins $93.42\pm 0.43\%$ $77.42\pm 4.92\%$
SwAV $93.87\pm 0.59\%$ $82.17\pm 0.05\%$
MoCo v3 $93.91\pm 0.60\%$ $78.82\pm 0.93\%$
ViT-DINO $90.53\pm 1.13\%$ $79.76\pm 2.15\%$
Semi-supervised (trained on labeled and unlabeled images) FixMatch $91.58\pm 0.65\%$ $83.34\pm 2.53\%$
@@
CLASS-M $\mathbf{95.35\pm 0.46\%}$ $\mathbf{92.13\pm 0.89\%}$
4.2 Adaptive stain separation
In experiments, we maintained slide-by-slide level adaptive stain separation based on the following considerations: The stain styles vary across different WSIs due to different conditions. But within one WSI, most conditions, such as stain manufacturer, storage condition, and staining procedure, remain consistent. On the other hand, if adaptive stain separation is applied patch-by-patch, there are not enough pixel samples inside each patch to obtain robust stain separation results. For example, in Figure 3(a), the top image in Necrosis category has very limited Hematoxylin portion, which makes stain separation less trustful at patch-by-patch level.
The RGB to H/E conversion matrices for each WSI were computed before the CLASS-M model training. We provide our original code for adaptive stain separation based on algorithms in  [40]. The link to our code can be found at github.com/BzhangURU/Paper_CLASS-M/tree/main/Section2_get_stain_separation_matrices
@@ 4.3 Experiment settings
For a fair comparison, we used ImageNet pre-trained ResNet18 model  [26] as an initialization for all Convolutional Neural Network (CNN) models in semi-supervised learning. We compared our CLASS-M model with other state-of-the-art semi-supervised learning models, including FixMatch  [50] and MixMatch  [3]. In CLASS-M model training, we performed jittering of brightness, contrast and saturation on original RGB images, as well as image augmentations on H images and E images after stain separation, such as random rotation, random crop to $256\times 256$, random flip and jittering of brightness. The image augmentations on H images and E images are independent, for example, they could have different rotation angles in random rotation. Since H images and E images are one-channel images, we summed the pre-trained weights on the first layer of ResNet18 as initialization weights.
@@
We used Root Mean Squared Propagation optimization method with a decaying learning rate. A batch size of 64 was chosen for all experiments. A balanced sampler was used to address the imbalance in the number of labels in the training set. In experiments with the Utah ccRCC dataset, we chose to have 32 labeled samples in each batch, with 8 samples for each category, while the remaining 32 samples were unlabeled. In experiments with the TCGA ccRCC dataset, we chose to have 33 labeled samples in each batch, with 11 samples for each category, and the remaining 31 samples would be unlabeled. In validation and test, we first calculated classification accuracies for each category, and then averaged them to get balanced validation accuracy and test accuracy. This approach ensures that each category holds equal importance, irrespective of the actual number of tiles for each category. The hyperparameters were fine-tuned for all models to get the best validation accuracy, then the test accuracy on that hyperparameter setting is reported as the final performance. We found that most hyperparameters didn’t need to be changed between Utah ccRCC dataset and TCGA ccRCC dataset. We chose decaying learning rate with initial value set to $10^{-4}$. We set the number of augmentation times K to 2, temperature $T$ to 0.5, $\alpha$ to 2, margin $m$ in contrastive loss to 37 and unlabeled $L_{2}$ loss weight to 7.5 for both datasets. We set contrastive loss weight to 0.1 for experiments on the Utah ccRCC dataset, and contrastive loss weight to 0.001 for the TCGA ccRCC dataset. For each epoch, we ran 1000 iterations and checked validation accuracy once. The total epochs were large enough to ensure convergence. The training was stopped if the best validation accuracy was no longer updated for more than 100 epochs. Each experiment was repeated three times to obtain the average test accuracy and standard deviation.
@@
The experiment platform is Python 3.7.11, Pytorch 1.9.0, torchvision 0.10.0, and CUDA 10.2. The GPUs we used are NVIDIA TITAN RTX. During CLASS-M model training, it usually took 13 GB GPU memory and 12 minutes for each training epoch. The code of our CLASS-M model is publicly available at github.com/BzhangURU/Paper_CLASS-M/tree/main/Section3_CLASS-M
@@
Table 2: Ablation studies of CLASS-M models on Utah ccRCC dataset and TCGA ccRCC dataset. Mean accuracy and standard deviation on test set are calculated. The results with best mean accuracy are shown in bold.
@@
CLASS-M $\mathbf{95.35\pm 0.46\%}$ $\mathbf{92.13\pm 0.89\%}$
CLASS-M without contrastive loss $90.92\pm 1.18\%$ $89.70\pm 0.77\%$
CLASS-M without aug on RGB images $94.41\pm 0.44\%$ $91.21\pm 2.05\%$
CLASS-M without adaptive stain separation (use fixed stain separation) $93.97\pm 0.40\%$ $90.97\pm 1.86\%$
@@
CLASS without adaptive stain separation (use fixed stain separation) $92.43\pm 0.34\%$ $79.89\pm 0.69\%$
@@
4.4 Comparison with baseline and other semi-supervised models
We tested our CLASS-M model along with other state-of-the-art semi-supervised models, including FixMatch and MixMatch using the Utah ccRCC dataset and TCGA ccRCC datasets. CLASS-M without pseudo-labeling using MixUp, which we name CLASS, was tested as a reference. The baseline ResNet18  [26] and Vision Transformer (ViT)  [17] were also tested where only labeled samples were used in training. The ResNet18 and ViT models were initialized with weights pre-trained on ImageNet.
As shown in Table 1, CLASS-M model outperforms not only the supervised baselines, but also other state-of-the-art models by a large margin on both datasets. We achieved a test accuracy of $95.35\pm 0.46\%$ on Utah ccRCC dataset, and $92.13\pm 0.89\%$ on TCGA ccRCC dataset. The ResNet18 baseline results from Table 1 indicate that TCGA ccRCC dataset presents a harder challenge. One possible reason is that the samples from Utah ccRCC dataset have less variation in staining procedure, and the storage conditions are uniform. However, on TCGA ccRCC dataset, slides come from different institutions, leading to more variation in the quality of slides. Additionally, the Necrosis tiles on TCGA training set only come from 4 WSIs, which causes more difficulties in learning robust features for this class. As shown in the recall section of Table 4 in Appenix B, pseudo-labeling with MixUp raised the Necrosis’s test accuracy from 60.47% to 86.65% on the TCGA ccRCC dataset, demonstrating that MixUp is especially effective on classes with very limited number of samples. In conclusion, our CLASS-M model achieves the highest accuracy results on a variety of tasks. We provide recall, precision and F-score for each run, with details available in Appendix B. After obtaining the CLASS-M models, we ran the models on WSIs from the test set and generated prediction heatmaps to better visualize patch-level predictions. The results can be found in Appendix C.
4.5 Comparison with self-supervised learning models
To compare with self-supervised learning approaches, we used both labeled and unlabeled samples to pre-train self-supervised learning models on the Utah ccRCC dataset and TCGA ccRCC dataset separately. After pre-training, the encoder was frozen, and a fully connected classification network was appended. The final fully connected classification layer was trained only on labeled training samples. We selected several state-of-the-art self-supervised learning models for our experiments, including Barlow Twins  [57], SwAV  [6], MoCo v3  [15], and ViT-DINO  [7], given their extensive use in histopathology [32]. As shown in Table 1, our proposed models, CLASS and CLASS-M outperform those self-supervised learning models on ccRCC datasets. The recall, precision and F-score of the self-supervised learning models can also be found in Appendix B. The disadvantage of the semi-supervised approach compared to self-supervised learning is the loss of generality for the encoder which must be retrained for each new task. However, our results demonstrate that a well-formulated semi-supervised learning approach can have accuracy advantages over a self-supervised pre-training approach due to end-to-end learning with labeled and unlabeled samples simultaneously. Therefore, if accuracy is of paramount importance, semi-supervised learning may be preferred over self-supervised learning.